A natural number x is chosen at random from the first 100 natural numbers.Then the probability,for the equation x+100/x>50 is:
(A)1/20
(B)11/20
(C)1/3
(D) none of these

Grade:12th pass

1 Answers

Arun
5 years ago
 

Total number of ways of selecting x is 100.

Now the given condition is x + 100/x > 50, on analyzing this equation, carefully, we see that this equation is satisfied for all the numbers x such that x > 48 and also for x = 1 and 2

So, favourable number cases is 55

Therefore, probability = 55/100 = 11/20